>Continuing Problem P12-47 Describing bonds, journalizing transactions for bonds payable using the straight-line amortizationc. Semiannual bond interest payments for 2020. Amortize the premium or discount using the straight-line amortization method.
Answer #1

1) Company received $208476 on issue of Bonds which has Face value $210000, hence the bond is issued at Discount amounting to $1524.

2)

a) Company has issued the Bonds and received $208476 which has Face Value of $210000.

Therefore, the company has received $208476 on Jan 2 2020.

b) Company has received $450000 on Jan 2 2020.

c) Semiannual Bond Interest Payment=($210000*7.5%)/2=$7875

Amortization of Discount of $1524 over 6 years using Straight Line Basis=$1524/6=$254

d) Payment on Mortgage payable=$45000

Interest Paid annually=$450000*8%=$36000

3) Total Interest Expense =$36000+$7875+$7875=$51750.